# CharacterGen: Efficient 3D Character Generation from Single Images with Multi-View Pose Canonicalization

### v2

@@ 3.3.1 Character Reconstruction with Coarse Texture
Inspired by LRM [14], we utilize a deep transformer network to efficiently reconstruct characters from the four-view images generated by the multi-view diffusion model. While LRM is trained using the Objaverse dataset [8] to allow versatile 3D object generation, it does not sufficiently capture the intricacies of human character layouts. To retain the reconstruction network’s ability to process both general 3D objects and stylized characters, we initially pre-train our transformer network on the Objaverse dataset. Then we fine-tune the model with our Anime3D dataset to introduce more priors of human body structure.
@@
Original LRM proposes to mainly train with NeRF [36] representation. However, directly extracting geometry from NeRF models often yields noisy surface geometry, which can be problematic for the subsequent use of character meshes in downstream graphics pipelines. Instead, we utilize a two-stage fine-tuning strategy for our reconstruction network. The first stage involves using a triplane NeRF representation similar to LRM, to establish the character’s coarse geometry and appearance. In the second stage, we modify the decoder module of our reconstruction network to predict signed distance functions (SDFs) rather than density fields, which allows CharacterGen to achieve smoother and more precise surface geometry.
@@
In addition to MSE loss, we also incorporate mask loss and LPIPS loss [75] to supervise reconstruction appearance. Mask loss is adopted on the rendering opacity to accelerate convergence and LPIPS loss is used to extract perceptual information from input images. The training target is given in Eq. 4.
$L_{\text{recon}}$ $=\lambda_{1}L_{\text{mse}}+\lambda_{2}L_{\text{mask}}+\lambda_{3}L_{\text{LPIPS}}$ (4)
Here, $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters, which are set to 1, 0.1 and 0.5 by default.
3.3.2 3D Character Refinement
@@
Our reconstruction network can rapidly reconstruct the 3D implicit representation of an character and we can extract the final mesh along with a coarse $UV$ map from the reconstructed tri-plane with DMTet [51]. However, the generated mesh still lacks texture details because the DMTet-based extraction process loses appearance information during the $UV$ unwrapping process. To tackle this problem, we further utilize the generated four-view images to improve the quality of the generated texture maps with texture back-projection techniques. For efficient rasterization in this step, we employ NvDiffRast [22] as the renderer. We find that directly optimizing all pixels of the generated texture map like NvDiffRec [40] with generated sparse input views leads to a texture down-sampling problem, resulting in severe degradation of the output refined texture map. To circumvent this problem, we project the four-view images into texels in texture space and employ a depth test to remove occluded texels. We also implement a normal-based filtering approach to remove pixels on the character’s silhouette to mitigate noisy texels. Then we utilize Poisson Blending [43] to aggregate projected texels and origin texels to reduce seams in the final textures.
4 Experiments
4.1 Data Preparation
For the training stage, we first divide our Anime3D dataset into a training set and a testing set in a 50:1 ratio. In the inference stage, we not only use our test set but also incorporate images from the Internet. Please refer to the supplementary materials for further implementation details of data processing and model training.
4.2 Results and Comparison
@@
We conduct experiments on both 2D multi-view character image generation and 3D character mesh generation to evaluate the efficiency of our CharacterGen. Additionally, we conduct user study on our generated characters and list the results in the appendix.
4.2.1 2D Multi-view Generation
@@
We test our models on images from the testing split of Anime3D as well as online sources and compare our results to those from Zero123 [30] and SyncDreamer [31]. The comparison results are shown in Fig. 8. It can be seen that, given some difficult body poses, Zero123 and SyncDreamer struggle to preserve enough geometry and appearance information of generated images. Our CharacterGen adeptly performs canonical pose calibration and generates consistent character images across four views, which significantly enhances the subsequent character mesh reconstruction process.
@@
We also conduct experiments on all character images from the test split of Anime3D and show the quality metrics in Tab. 1. Our calibrated A-pose images are benchmarked against ground-truth A-pose images, whereas images generated by other methods are compared with corresponding posed images. The results evaluate CharacterGen’s superior generation quality and its consistency with the multi-view diffusion model.
@@
Table 1: We show the quantitative metrics of all methods on the test split to evaluate the efficiency of our CharacterGen .
@@
CharacterGen 0.833 0.087 0.017
Zero123 [30] 0.768 0.224 0.142
SyncDreamer [31] 0.807 0.194 0.396
4.2.2 3D Character Generation
In this section, we compare our generated 3D characters with image-prompt 3D character generation methods. ImageDream [60] and Magic123 [46] all utilize the SDS-based optimization. TeCH [17] extends ECON [67] and utilizes DreamBooth [48] to achieve image-prompt generation. We visualize the results in Fig. 9.
It can be observed that our CharacterGen effectively avoid Janus problem thanks to our robust four-view reconstruction mechanism. Our generated 3D character meshes also exhibit satisfactory appearance for unseen body parts, with resourceful back-view and side-view priors from our Anime3D. Most 3D characters generated by other methods suffer from several mesh face cohesion problems, which makes it extremely hard to rig and animate these characters. CharacterGen can successfully generate canonical pose meshes from characters with tricky poses, which facilitates downstream graphic applications. We also evaluate other methods using A-pose character images canonicalized by Animate Anyone [15]. Please refer to the appendix for further details.
4.2.3 Generation Speed
Table 2: Time to generate a single 3D character. Models loading time is excluded for all methods.
@@ 4.2.3 Generation Speed
Magic123 [46] 70min
ImageDream [60] 45min
TeCH [17] 270min
We compare the time needed to generate a single 3D character by our method and other image-prompted 3D generation methods, with the comparison results detailed in Tab. 2. Our method is significantly faster than other alternatives.
4.3 Ablation Study
4.3.1 IDUNet
Figure 5: Frozen IDUNet cannot extract sufficient appearance information from the prompt image and generates dissimilar images.
@@ 4.3.1 IDUNet
4.3.2 Pose Embedding Network
@@
Figure 6: Without the pose embedding network, the generated characters may be misplaced.
4.4 Applications
CharacterGen can generate A-pose 3D characters with detailed texture maps, thereby simplifying the subsequent rigging process. We employ AccuRig [1] to automatically rig our generated character meshes. The rigged 3D characters can be readily utilized as animated 3D assets in various domains. We render various animated rigged models in Warudo [11] and present some results in Fig. 7.
Figure 7: We rig the generated characters and utilize them as 3D assets in downstream applications.
5 Limitations and Discussion
While our method can generate 3D characters from a single input image in a arbitrary pose, certain limitations still exist. For the four-view A-pose image generation step, our method may not retain enough information when the character is in an extreme pose or is rendered from an non-common viewpoint.
@@ 6 Conclusions
Figure 8: We compare our generated four A-pose character images with other methods. The azimuths for all examples are set as $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$.
Figure 9: We compare the appearance and geometry of our generated 3D characters with other methods.
Supplementary Material
7 Implementation Details
For our Anime3D dataset, we utilize three-vrm [44] to set character poses. For “A-pose” characters, we set the joint rotation of the left and right arm to $45^{\circ}$ in the $Z$-axis and the left and right upper leg to $6^{\circ}$ in the $Z$-axis. All other joint parameters remain untouched. For the posed character setting, we download 10 human skeleton animations from Mixamo [37], including sitting, singing, and walking, etc. We randomly select frames from these animations and apply the corresponding motion to the VRM character models. In addition, we also randomize joints of the mouth and eyes to generate a variety of facial expressions, such as winking. For the dataset rendering settings. We configure the camera field of view (FoV) to $40^{\circ}$. We also set the distance between the camera and the scene origin to 1.5m.
We use the Stable Diffusion 2.1 model as the base model of both our IDUNet and Multi-view UNet. The training process is carried out on 8 NVIDIA A100 GPUs for 3 days on images with $512\times 512$ resolution and another 2 days on images with $768\times 512$ resolution. In each training step, we jointly train both IDUNet and Multi-view UNet. We begin by sampling a pair of four-view images and a single-condition posed image from Anime3D. Given our intention to generate four images with azimuths of $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$, we sample this group with a probability of 0.8. We also include other four-view image groups in our training steps to strengthen the diffusion model’s spatial understanding of 3D characters. We sample the front-view posed character images whose azimuth ranges from $90^{\circ},90^{\circ}]$. During the training phase, the posed image is fed into the IDUNet, and the four-view images are sent to the Multi-view UNet
To fine-tune the transformer-based reconstruction model, we first fine-tune for 50 epochs with the NeRF representation and for another 30 epochs on SDF. The fine-tuning process takes 1 day on 8 A100 GPUs. As CharacterGen does not require training in the inference steps, the entire generation pipeline (including four canonical-pose image generation, 3D character mesh reconstruction, and texture map refinement) can run on a single GPU.
Figure 10: We show the appearance and geometry of other methods with A-posed input generated by Animate Anyone and compare our results with them.
8 Comparison with A-pose character generation
We also utilize concurrent work Animate Anyone [15] to achieve the pose canonicalization process and conduct experiments on other competing 3D generation methods with these A-pose images. Since the official implementation of Animate Anyone is not open-sourced, we adopt MooreThreads [38]’s implementation. Results are shown in Fig. 10. Compared to our results, it can be observed that our methods can preserve more geometric and appearance features with our multi-view diffusion models, showing the effectiveness of our pose canonicalization model. Furthermore, the use of four-view A-pose images provides more comprehensive character features for the reconstruction network in the subsequent stage. This leads to a smoother geometry and more detailed texture in our reconstructed 3D characters.
9 User Study
To better evaluate the robust generation ability of our CharacterGen, we collect 15 generated four-view character images and 10 character meshes for our user study. We compare results generated by CharacterGen with other methods mentioned in the experiment section. We ask 21 volunteers to first evaluate the style consistency between the condition images and the generated four-view images, as well as the spatial consistency within the generated multi-view images. Then, they are required to assess the geometry quality and texture quality of generated 3D character models. For each example, the volunteers are asked to vote for the one they consider to be the best. As shown in Fig 11, our CharacterGen receives a significantly higher preference compared to methods for both 2D and 3D generation tasks.
10 Realistic Character Generation
Figure 11: Statistics of the user study.
Figure 12: Generated results of realistic humans.
Figure 13: We compare the appearance and geometry of our generated 3D characters with other methods using A-pose input images canonicalized by Animate Anyone.
We also test our CharacterGen on some realistic posed human images and demonstrate an example in Fig. 12. It can be observed that CharacterGen can successfully generate high-consistent 3D realistic characters even though it is only trained on stylized character data. Combining additional realistic character data in the future training stage may further bridge the domain gap and produce realistic characters with higher quality.
11 A-pose rigging
To better evaluate how using an A-pose character aids skeleton rigging and animation process, we rig two 3D character meshes generated by our CharacterGen and ImageDream [60] and visualize the animated characters in Fig. 13. It can be observed that non-A-pose characters result in a severe mesh cohesion problem and the body structure is severely distorted while our A-posed character can be successfully animated.
12 Additional Quantitative Evaluation
Table 3: We utilize CLIP score to assess style similarity between generated character and given condition images for both 2D and 3D tasks.
method CLIP score (multi-view 2D) method CLIP score (3D rendered)
CharacterGen $\mathbf{83.69}$ CharacterGen $\mathbf{79.77}$
Zero123 [30] 79.02 Magic123 [46] 74.71
SyncDreamer [31] 75.41 ImageDream [60] 73.65
We additionally quantitatively evaluate the coherence between our generated results and the input images, using the same candidates as in the user study. We adopt the CLIP score as the metric and utilize ViT-B/32 as the backbone model. For 2D multi-view image generation, we calculate the CLIP scores of the input image and four generated images. For 3D character generation, we first render the 3D representation at an azimuth interval of $3^{\circ}$, and then compute the CLIP scores between the input image and all the rendered frames. As displayed in Tab. 3, our CharacterGen gains superior results in both 2D and 3D generation, evaluating the robust appearance modeling ability of our IDUNet, multi-view UNet, and the subsequent reconstruction stage.

### v3

@@ 3.3.1. Character Reconstruction with Coarse Texture
Inspired by LRM (Hong et al., 2023b), we utilize a deep transformer network to efficiently reconstruct characters from the four-view images generated by the multi-view diffusion model. While LRM is trained using the Objaverse dataset (Deitke et al., 2023) to allow versatile 3D object generation, it does not sufficiently capture the intricacies of human character layouts. To retain the reconstruction network’s ability to process both general 3D objects and stylized characters, we initially pre-train our transformer network on the Objaverse dataset. Then we fine-tune the model with our Anime3D dataset to introduce more priors of human body structure.
@@
Original LRM proposes to mainly train with NeRF (Mildenhall et al., 2022) representation. However, directly extracting geometry from NeRF models often yields noisy surface geometry, which can be problematic for the subsequent use of character meshes in downstream graphics pipelines. Instead, we utilize a two-stage fine-tuning strategy for our reconstruction network. The first stage involves using a triplane NeRF representation similar to LRM, to establish the character’s coarse geometry and appearance. In the second stage, we modify the decoder module of our reconstruction network to predict signed distance functions (SDFs) rather than density fields, which allows CharacterGen to achieve smoother and more precise surface geometry.
@@
In addition to MSE loss, we also incorporate mask loss and LPIPS loss (Zhang et al., 2018) to supervise reconstruction appearance. We adopt the binary-cross-entropy loss between ground-truth alpha masks and rendered alpha masks as the mask loss to help the reconstruction model distinguish empty space within input four-view images. LPIPS loss is used to extract perceptual information from input images. The training target is given in Eq. 4.
(4) $L_{\text{recon}}$ $=\lambda_{1}L_{\text{mse}}+\lambda_{2}L_{\text{mask}}+\lambda_{3}L_{\text{LPIPS}}$
Here, $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters, which are set to 1, 0.1 and 0.5 by default. We leverage Laplacian Smooth on the extracted meshes to further reduce the noise on the surface.
3.3.2. 3D Character Refinement
@@
Our reconstruction network can rapidly reconstruct the 3D implicit representation of a character and we can extract the final mesh along with a coarse $UV$ map from the reconstructed tri-plane with DMTet (Shen et al., 2021). However, the generated mesh still lacks texture details because the DMTet-based extraction process loses appearance information during the $UV$ unwrapping process. To tackle this problem, we further utilize the generated four-view images to improve the quality of the generated texture maps. For efficient rasterization in this step, we employ NvDiffRast (Laine et al., 2020) as the renderer. Since the generated four-view images are in lower resolution than the texture map, multiple texels may be projected onto the same image pixel. During differentiable-rendering-based optimization, the gradients for these texels become noisy. The sparsity of input views compounds the difficulty of fixing these noisy texels, resulting in severe degradation of the output refined texture map. To circumvent this problem, we project the four-view images into texels in texture space and employ a depth test to remove occluded texels. We also notice that directly back-projecting the four-view images onto the mesh leads to noisy texels at the character’s body silhouette. We compute the inner products of the four orthogonal camera view directions with the normal texture map. Texels with inner products greater than -0.2 are disregarded to eliminate noise around the silhouette. For texels overlapping in multiple views, we select the back-projected texels with RGB values closest to the coarse texture. Then we utilize Poisson Blending (Pérez et al., 2003) to aggregate projected texels and origin texels to reduce seams in the final textures.
4. Experiments
4.1. Implementation Details
We divide our Anime3D dataset into a training set and a testing set in a 50:1 ratio. We use the Stable Diffusion 2.1 model as the base model of both our IDUNet and multi-view UNet. The training process is carried out on 8 NVIDIA A800 GPUs for 3 days on images with $512\times 512$ resolution and another 2 days on images with $768\times 512$ resolution. In each training step, we jointly train both IDUNet and multi-view UNet. We begin by sampling a group of four-view images and a single-condition posed image from Anime3D. Given our intention to generate four images with azimuths of $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$, we sample this group with a probability of 0.8. We also include other four-view image groups in our training steps to strengthen the diffusion model’s spatial understanding of 3D characters. We sample the front-view posed character images whose azimuth ranges from $[-90^{\circ},90^{\circ}]$. During the training phase, the posed image is fed into the IDUNet, and the four-view images are sent to the multi-view UNet.
To fine-tune the transformer-based reconstruction model, we first fine-tune for 50 epochs with the NeRF representation and for another 30 epochs on SDF. The fine-tuning process takes 1 day on 8 A800 GPUs. As CharacterGen does not require training in the inference steps, the entire generation pipeline (including four canonical-pose image generation, 3D character mesh reconstruction, and texture map refinement) can run on a single GPU.
4.2. Results and Comparison
@@
We conduct experiments on both 2D multi-view character image generation and 3D character mesh generation to evaluate the efficiency and effectiveness of our CharacterGen.
4.2.1. 2D Multi-view Generation
@@
We test our models on images from the testing split of Anime3D as well as online sources and compare our results to those from Zero123 (Liu et al., 2023c) and SyncDreamer (Liu et al., 2023a). The comparison results are shown in Fig. 5. It can be seen that, given some difficult body poses, Zero123 and SyncDreamer struggle to preserve enough geometry and appearance information of generated images. Our CharacterGen adeptly performs canonical pose calibration and generates consistent character images across four views, which significantly enhances the subsequent character mesh reconstruction process.
@@
We also conduct experiments on all character images from the test split of Anime3D. We fine-tune the baseline methods on Anime3D for 100 epochs and show the quality metrics in the upper part of Tab. 1. Our calibrated A-pose images are benchmarked against ground-truth A-pose images, whereas images generated by other methods are compared with corresponding posed images. The results evaluate CharacterGen’s superior generation quality and its consistency with the multi-view diffusion model.
@@
Table 1. We show the quantitative metrics of both 2D multi-view generation and 3D character generation methods on the test split of Anime3D to evaluate the effectiveness of our CharacterGen.
@@
CharacterGen(2D) 0.901 0.086 0.019 -
Zero123 0.768 0.224 1.42 -
Zero123(fine-tuned) 0.813 0.175 1.34 -
SyncDreamer 0.807 0.194 0.396 -
SyncDreamer(fine-tuned) 0.822 0.17 0.37 -
IP-Adapter+SDXL 0.845 0.143 0.074 -
CharacterGen(3D) 0.898 0.093 0.032 0.001
Magic123 0.873 0.134 0.116 0.0034
ImageDream 0.886 0.11 0.345 0.002
[…]
4.2.2. 3D Character Generation
In this section, we compare our generated 3D characters with image-prompt 3D character generation methods. ImageDream (Wang and Shi, 2023) and Magic123 (Qian et al., 2023) all utilize the SDS-based optimization. TeCH (Huang et al., 2023b) extends ECON (Xiu et al., 2023) and utilizes DreamBooth (Ruiz et al., 2023) to achieve image-prompt generation. We visualize the results in Fig. 6. We also conduct quantitative experiments on the 3D generation results on the test split of Anime3D. The texture quality metrics are obtained by comparing rendered images and ground-truth images across the four orthogonal views. We choose Chamfer Distance(CD) as the metric to evaluate the geometry quality of generated meshes. We normalize the meshes to $[-0.5,0.5]^{3}$ for calculating CD. The quantitative results are listed in the lower part of Tab. 1.
It can be observed that our CharacterGen effectively avoids “Janus” problem thanks to our robust four-view reconstruction mechanism. Our generated 3D character meshes also exhibit satisfactory appearance for unseen body parts, with resourceful back-view and side-view priors from our Anime3D. Most 3D characters generated by other methods suffer from several mesh face cohesion problems, which makes it extremely hard to rig and animate these characters. CharacterGen can successfully generate canonical pose meshes from characters with tricky poses, which facilitates downstream graphic applications. We also evaluate other methods using A-pose character images canonicalized by Animate Anyone (Hu et al., 2023). Please refer to the supplementary materials for further details.
4.2.3. Comparison with IP-Adapter
Previous work IP-Adapter (Ye et al., 2023) also supports the image-prompt generation task by incorporating adapter modules into diffusion models. We notice that IP-Adapter does not include pre-trained models for Stable Diffusion 2.1, which is the base model of our multi-view UNet. Alternatively, we train an SDXL-base model to generate $2\times 2$ grid character images in four views with the azimuth of $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ for 100 epochs. The grid images are organized with a resolution of $1024\times 1024$, which is the standard resolution of SDXL. Then we integrate official pre-trained IP-Adapter-SDXL into the base SDXL model to obtain image-conditioned multi-view results. We show the visualization results in Fig. 7 and quantitative results in Tab. 1. It can be observed that IP-Adapter cannot preserve detailed appearance information and may fail to generate correct character layout, while CharacterGen can effectively generate high-consistent multi-view character images with our IDUNet.
Figure 7. We compare the results of CharacterGen and IP-Adapter-SDXL.
4.2.4. Generation Speed
Table 2. Time to generate a single 3D character. Models loading time is excluded for all methods.
@@ 4.2.4. Generation Speed
Magic123 (Qian et al., 2023) 70min
ImageDream (Wang and Shi, 2023) 45min
TeCH (Huang et al., 2023b) 270min
We compare the time needed to generate a single 3D character mesh by our method and other image-prompted 3D generation methods, with the comparison results detailed in Tab. 2. Our method is significantly faster than other alternatives. SyncDreamer and Zero123 are used to generate multi-view images. Their time cost varies based on the chosen 3D reconstruction methods. The default NeuS (Wang et al., 2021a) reconstruction takes about 10 minutes.
4.3. User Study
Table 3. Statistics of the user study. We display the voting results for both 2D multi-view images and 3D textured character meshes.
metric CharacterGen Zero123(2D) SyncDreamer(2D) Magic123(3D) ImageDream(3D)
2D multi-view style consistency 85.4% 10.5% 4.1% - -
2D multi-view consistency 81.0% 17.1% 1.9% - -
3D character geometry quality 78.6% - - 2.86% 18.6%
3D character texture quality 87.1% - - 1.9% 11.0%
To better evaluate the robust generation ability of our CharacterGen, we collect 15 generated four-view character images and 10 character meshes for our user study. We compare results generated by CharacterGen with other methods mentioned in Sec. 4.2. We ask 21 volunteers to first evaluate the style consistency between the condition images and the generated four-view images, as well as the spatial consistency within the generated multi-view images. Then, they are required to assess the geometry quality and texture quality of generated 3D character models. For each example, the volunteers are asked to vote for the one they consider to be the best. As shown in Tab 3, our CharacterGen receives a significantly higher preference compared to other methods for both 2D and 3D generation tasks.
Table 4. We utilize CLIP score to assess style similarity between generated characters and given condition images with 2D multi-view images and 3D rendered images.
method(2D) CLIP score method(3D) CLIP score
CharacterGen $\mathbf{83.69}$ CharacterGen $\mathbf{79.77}$
Zero123 79.02 Magic123 74.71
SyncDreamer 75.41 ImageDream 73.65
We additionally quantitatively evaluate the coherence between our generated results and the input images the same as in the user study. We adopt the CLIP score as the metric and utilize ViT-B/32 as the backbone model. For 2D multi-view image generation, we calculate the CLIP scores of the input image and four generated images. For 3D character generation, we first render the 3D representation at an azimuth interval of $3^{\circ}$, and then compute the CLIP scores between the input image and all the rendered frames. As displayed in Tab. 4, our CharacterGen gains superior results in both 2D and 3D generation, evaluating the robust appearance modeling ability of our IDUNet, multi-view UNet, and the subsequent reconstruction model.
4.4. Ablation Study
4.4.1. IDUNet
@@
Figure 8. Frozen IDUNet cannot extract sufficient appearance information from the prompt image and generates dissimilar images.
4.4.2. Pose Embedding Network
@@
Figure 9. Without the pose embedding network, the generated characters may be misplaced.
4.5. Applications
CharacterGen can generate A-pose 3D characters with detailed texture maps, thereby simplifying the subsequent rigging process. We employ AccuRig (actorcore, 2023) to automatically rig our generated character meshes. The rigged 3D characters can be readily utilized as animated 3D assets in various domains. We render various animated rigged models in Warudo (HakuyaLabs, 2023) and present some results in Fig. 10.
Figure 10. We rig the generated characters and utilize them as 3D assets in downstream applications.
To better evaluate how using an A-pose character aids skeleton rigging and animation process, we rig two 3D character meshes generated by our CharacterGen and ImageDream (Wang and Shi, 2023) and visualize the animated characters in Fig. 11. It can be observed that non-A-pose characters encounter severe mesh cohesion problem and the body structure is severely distorted while our A-posed character can be successfully animated.
Figure 11. We compare CharacterGen’s animated 3D characters with ImageDream (Wang and Shi, 2023)
5. Limitations and Discussion
While our method can generate 3D characters from a single input image in an arbitrary pose, certain limitations still exist. For the four-view A-pose image generation step, our method may not retain enough information when the character is in an extreme pose or is rendered from a non-common viewpoint.
@@ 6. Conclusions
Acknowledgements.
This work is supported by the National Science and Technology Major Project (2021ZD0112902), the National Natural Science Foundation of China (Grant No. 62220106003), and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology. The authors would like to thank Ralph R. Martin, Yuan-Chen Guo, and Yang-Guang Li for helpful discussion.